OpenStudy (anonymous):

Which fractions are equivalent to 25/45 A. 5/9,15/28 B.30/48,75/135 C.5/9,75/135 D.35/54,40/108

OpenStudy (thread252):

The answer should be C 25/45=75/135=5/9

OpenStudy (thread252):

if i get it correct an i have a medal? :) please?